<p>14:05</p>
<p>Run 22 stopped. Run 23 started we will now adjust the vertical beam position to see if we observe the beam stripe seen in the last beamtime.</p>
<p>Current vertical beam position 0.0 mm. Horizontal beam position 7.0 mm. Coasting beam.</p>
<p>Beam D+, 1.5 MeV. Storage time ~ 20 minutes. Target ON. Density ~1e11 at/cm2.</p>
<p>XY plot from run 22 with the same settings - attachment 1</p>
<p>Beam stripe looks like it is just on bottom detectors (just )</p>
<p>14:10</p>
<p>New beam settings</p>
<p>Vertical Beam position 5.0mm, Horizontal unchanged. Storage time reduced to ~2 minute</p>
<p>No real change observed. Perhaps some thing more can be seen on top right detector but definitely not like in the Feb experiment.</p>
<p>14:15</p>
<p>New beam settings</p>
<p>Vertical Beam position 9.0mm, Horizontal unchanged.</p>
<p>No real change observed again.</p>
<p>14:20</p>
<p>New beam settings</p>
<p>Vertical Beam position -9.0mm, Horizontal unchanged.</p>
<p>More events observed on bottom left detector but nothing like the Feb measurements. I think we do not have stripe effects like previously.</p>
<p>14:40:50</p>
<p>New Beam settings to check beam lifetime.</p>
<p>Vertical beam position set back to 0.0 mm. Horizontal unchanged</p>
<p>In bunched mode to observed intensity over time. </p>
<p>Injection into ring 2.2e7</p>
<p>15:15</p>
<p>Beam intensity dropped to 1.1e7. Beam lifetime 35 minutes.</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p> </p>
<p>15:24</p>
<p>We removed the target. Beam intensity at 2.2E7</p>
<p>After 35 minutes, the beam intensity is 1.41e7. So target does have effect on beam lifetime! We would like more.</p>
<p>16:05</p>
<p>We will now bring target back in and try to increase density.</p>
<p>16:10</p>
<p>Some sort of GSI-wide failure, we have lost beam entirely. We don't see any effect on our system. Temperatures and leakage currents are stable</p>
<p>Detectors moved fully out for safety</p>
<p>17:15</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>19:15</p>
<p>Still no beam</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>21:15</p>
<p>Still no beam</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>21:40</p>
<p>Beam has returned!</p>
<p>22:00</p>
<p>Target at around 9e11 and appears stable</p>
<p>Merger OK, Tape server OK</p>
<p>ADC data item statistics looks as before. OK</p>
<p>FEE temperatures look OK</p>
<p>System checks OK</p>
<p>Leakage current stable</p>
<p>22:33</p>
<p>Beam back in the ring. Beam settings to be updated.</p>
<p>Rates in right hand detectors look slightly different to before - lower in bottom right, slightly higher in top left. Perhaps beam vertical position needs to be lowered slightly.</p>
<p>ASIC check load just to be sure. Rates in aida15 greatly reduced - only temporarily.</p>
<p> </p> |